The width of a box is 3 feet less than the length. The height is 4 feet less than the length. The volume of the box is 36 cubic feet. Find the length of box .
Answer by Gogonati(855)

Solution:length x feet; Width x-3 feet; Height x-4 feet. Since the volume is 36 cubic feet we write: x(x-3)(x-4)=36; (x^2-3x)(x-4)=36; x^3-4x^2-3x^2+12=36;x=6
Answer:The length of the box is 6 feet.